OpenStudy (anonymous):

Evaluate expression : C 6 2 _____ = C 6 1

OpenStudy (anonymous):

i am going to guess it is \[\frac{_6C_2}{_6C_1}\]

OpenStudy (anonymous):

\[_6C_1=6\]since there are six ways to choose one thing out of six

OpenStudy (anonymous):

\[_6C_2=\frac{6\times 5}{2}=15\]

OpenStudy (anonymous):

Thank you, I am still so confused am I making this harder than it really is??

OpenStudy (anonymous):

Is there any way to learn these that you may think could help?

OpenStudy (anonymous):

I have been struggling with these for 2 days now